Grade 12th passPhysical Chemistry

A litre contains 2 mole of PCl5 initially. If Kc is found to be 1, the degree of dissociation of PCl5 is?

Abhinandan Kumar madheshiya
7 Years agoGrade 12th pass

1 Answer

Vikas TU
7 Years ago
PCl5 dissociates to give PCl3 and Cl2 as it is generally unstable because of the butterfly structure it comes with.  Also it is given that Kc is 1.
So the reaction is PCl5 is in equilibrium with PCl3 and Cl2.
                             C                                         0             0
                         C-Ca                                    Ca             Ca
Where ‘a’ is the dissociation constant.
Since Kc is 1 (Ca)^2=C-Ca;
This on further simplification and assuming that 1-a is approximately 1 we get a to be
a=(Kc/C)^½= 0.717.